OpenStudy (anonymous):

solve each quadratic equation. show your work. 1. (2x-1)(x+7)=0 2. x^2+3x=10 3. 4x^2=100

OpenStudy (anonymous):

1. x=12,−7 2. x^2+3x=10 x=±5

OpenStudy (anonymous):

For equation 1: Ask: When will (2x−1)(x+7) equal zero? Answer: When 2x−1=0 or x+7=0. Solve each of the 2 equations above:

OpenStudy (campbell_st):

in 1. if either factor equals zero then the quadratic will equal zero. so you need to find the values of x that make 2x - 1 = 0 and x + 7 = 0 you will get 2 different values 2. rewrite the equation as \[x^2 + 3x - 10 = 0\] factor the equation then solve the factors just like question 1. 3. a lot simpler... divide both sides of the equation by 4 then take the square root of both sides... remember there are 2 answers, a positive and negative hope it helps

OpenStudy (anonymous):

For equation 2: x2+3x−10=0 x2+3x−10 (x−2)(x+5)=0 x=2,−5

OpenStudy (campbell_st):

there you go... someone always posts answers, even though Open Study asks that you don't